OpenStudy (anonymous):

Which is a factor of 8x^2 - 38x - 10

OpenStudy (anonymous):

4 x+1 x-5 3x+1

OpenStudy (anonymous):

i know you have to find the GCF which is 2

OpenStudy (anonymous):

2( 4x^2 - 19x - 5)

OpenStudy (swissgirl):

yup

OpenStudy (swissgirl):

ok now we need to find two numbers that when you multiply them it equals to a*c and when you add them it equals to b Our equation is in the form of \(ax^2+bx+c \) so our a=4 b=-19 and c=-5 \(a*c=4*(-5)=-20 \) So what are two numbers that when you multiply them it equals -20 and when you add them it equals -19?

OpenStudy (anonymous):

thats what i was confused about :-/

OpenStudy (swissgirl):

lol So are u fine from here? Can u continue the rest on ur own

OpenStudy (anonymous):

almost....

OpenStudy (swissgirl):

ok so what are the 2 numbers firstly?

OpenStudy (anonymous):

sooooo it has to = 20x^2

OpenStudy (swissgirl):

what has to equal \( 20x^2 \)?

OpenStudy (swissgirl):

ok the 2 numbers when you multiply them have to equal -20 and when u add them it has to equal -19

OpenStudy (anonymous):

lol i have no clue what two numbers add up to 19 and multiply to 20x^2

OpenStudy (swissgirl):

not 20x^2 its just simply -20

OpenStudy (swissgirl):

So what are two numbers that when you multiply them it equals -20 and when you add them it equals -19?

OpenStudy (anonymous):

when you find the GCF which is 4 it would be 4(2x^2 -19x -5)

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

i ment 2 ( 4x^2 - 19 -5)

OpenStudy (anonymous):

so wouldnt it be 20x^2

OpenStudy (anonymous):

2(x-5)(4x+1)

OpenStudy (anonymous):

ohhh i understand now....

OpenStudy (anonymous):

What are your options? 2 x-5 or 4x+1 on that list?

OpenStudy (swissgirl):

Sorry abt that. My bro needed my comp emergency but ill just write out the steps neways

OpenStudy (swissgirl):

So the two numbers are 1and -20 1+(-20)=-19 1*(-20)=-20 so we now replace the -19x with 1x and -20x \(2(4x^2+1x-20x-5)=0 \) \( 2(x(4x+1)-5(4x+1))=0 \) \( 2(x-5)(4x+1)=0 \) so as we see (x-5) is the answer